# MorphArtGrasp: Morphology-Aware Cross-Embodiment Dexterous Hand Articulation Generation for Grasping

### v2

@@ Abstract
Dexterous grasping with multi-fingered hands remains challenging due to high-dimensional articulations and the cost of optimization-based pipelines. Existing end-to-end methods require training on large-scale datasets for specific hands, limiting their ability to generalize across different embodiments. We propose MorphArtGrasp, an eigengrasp-based, end-to-end framework for cross-embodiment grasp generation. From a hand’s morphology description, we derive a morphology embedding and an eigengrasp set. Conditioned on these, together with the object point cloud and wrist pose, an amplitude predictor regresses articulation coefficients in a low-dimensional space, which are decoded into full joint articulations. Articulation learning is supervised with a Kinematic-Aware Articulation Loss (KAL) that emphasizes fingertip-relevant motions and injects morphology-specific structure. In simulation on unseen objects across three dexterous hands, MorphArtGrasp attains a 91.9% average grasp success rate with $\mathord{<}0.4\,\mathrm{s}$ inference per grasp. With few-shot adaptation to an unseen hand, it achieves 85.6% success on unseen objects in simulation, and real-world experiments on this few-shot-generalized hand achieve an 87% success rate. The code and additional materials are available on our project website https://connor-zh.github.io/MorphArtGrasp/.
@@ I INTRODUCTION
In this paper, we propose MorphArtGrasp, a framework for cross-embodiment dexterous grasp generation. MorphArtGrasp extracts eigengrasps and morphology embeddings directly from a robotic hand’s Unified Robot Description Format (URDF), explicitly incorporating kinematic and geometric constraints. Given an object point cloud and a target wrist pose, an amplitude predictor estimates articulation coefficients in the eigengrasp space, which are then decoded into full joint configurations. To supervise training, we introduce a Kinematic-Aware Articulation Loss (KAL) that emphasizes fingertip-relevant motions and implicitly encodes morphology-specific information, in contrast to naïve per-joint regression objectives (e.g., mean squared error) that overemphasize raw joint deviations.
We demonstrate the effectiveness of the proposed framework through extensive evaluations in both simulation and real-world settings. In simulation on unseen objects across three dexterous hands, MorphArtGrasp achieves a 91.9% average grasp success rate with $0.4\,\text{s}$ inference per grasp; with few-shot adaptation to an unseen hand it attains 85.6% success on unseen objects in simulation, and real-world experiments on the same few-shot–generalized hand achieve 87% success.
@@
  • We propose MorphArtGrasp, an eigengrasp-based framework for cross-embodiment dexterous grasp generation that predicts articulations in an end-to-end manner.
@@ II-B Cross-Embodiment Dexterous Grasping
Another line of research focuses on enabling dexterous grasping across different robotic hands by leveraging hand-agnostic representations or shared action spaces. Cross-embodiment reinforcement learning approaches, such as the eigengrasp-based method by Yuan et al. [24], employ a shared low-dimensional action space to allow a single policy to operate across multiple hands, although training is often sample-intensive. More recently, GET-Zero [12] introduces a graph-based embodiment transformer that conditions policy learning on the structural graph of the hand, achieving zero-shot transfer to unseen morphologies in dexterous in-hand manipulation tasks. While highly effective for manipulation, its focus is on policy learning rather than generating feasible grasps, leaving a gap that MorphArtGrasp addresses by focusing on effective grasp generation for diverse hand morphologies.
@@
Figure 2: Method Overview. MorphArtGrasp processes hand URDF, object point cloud, and wrist pose through specialized encoders to generate morphology embedding $\boldsymbol{m}$, point cloud embedding $\boldsymbol{p}$, and pose encoding $\boldsymbol{h}$. These are fused to predict eigengrasp amplitudes $\boldsymbol{a}$, which are combined with eigengrasps $E$ to produce the final articulation vector $\boldsymbol{q}$.
@@ III-B Method Overview
An overview of MorphArtGrasp is shown in Fig. 2. The framework takes as input the hand URDF, the object point cloud, and the wrist pose. The hand URDF is first processed by the morphology encoder (Sec. III-C), which produces a morphology embedding ($\boldsymbol{m}$) together with a set of eigengrasps ($E$). The object point cloud is then processed by the object encoder (Sec. III-D) to extract a point cloud embedding ($\boldsymbol{p}$). These embeddings, combined with the wrist pose ($\boldsymbol{h}$), are fused to form conditioned eigengrasp tokens. The amplitude predictor (Sec. III-E) takes these tokens as input and estimates the amplitudes for each eigengrasp ($a$). Finally, the articulation vector ($\boldsymbol{q}$) is obtained by the dot product between the predicted amplitudes and the eigengrasps, and together with the wrist pose, defines the final hand pose.
@@ III-C Morphology Encoder
A central component of MorphArtGrasp is the morphology encoder (illustrated in Fig. 3), which learns structured representations of dexterous hands directly from their URDF descriptions. In contrast to prior work that encodes hand morphology solely as mesh point clouds [15, 9, 20], MorphArtGrasp leverages the URDF to extract explicit kinematic constraints and geometric primitives. Based on this structured information, the encoder predicts both the eigengrasps of the hand and a compact latent representation that captures its underlying morphology.
@@ IV EXPERIMENTS
  • Q1: How well does MorphArtGrasp perform in realistic settings where wrist poses are predicted by a naive pose predictor, compared to state-of-the-art baselines?
  • Q2: Does integrating articulation predictions from MorphArtGrasp with the wrist poses generated by existing baselines (GraspIt, DexGraspNet, DRO) improve their grasp success rates?
  • Q3: Can MorphArtGrasp generalize to previously unseen robotic hands with only few-shot adaptation?
@@ IV-A Datasets
We train MorphArtGrasp on three dexterous robotic hands: ShadowHand (5 fingers, 22 DoF), Allegro Hand (4 fingers, 16 DoF), and Barrett Hand (3 fingers, 8 DoF). For Allegro and Barrett, we adopt grasp data from the MultiGripperGrasp dataset [2]. Since the ShadowHand subset in MultiGripperGrasp suffers from severe self-collision artifacts, we instead use DexGraspNet [20] for ShadowHand data.
@@ IV-C Articulation Prediction with Predicted Wrist Pose
To answer Q1, we evaluate MorphArtGrasp in a realistic setting where wrist poses are predicted rather than provided. Specifically, we use 6-DoF GraspNet [11] to generate parallel-gripper candidates per object, and map each candidate to a hand-specific wrist pose using a calibrated rigid transform $\Delta_{\text{hand}}$:
@@
For each retained wrist pose $T_{\text{wrist}}^{(\text{hand})}$, MorphArtGrasp predicts an articulation, and the resulting grasp is evaluated in simulation. Experiments are conducted on a test set of 28 unseen objects, with 50 candidate grasps per object. Example grasps generated by MorphArtGrasp are visualized in Fig. 1.
@@
As an ablation, we also compare MorphArtGrasp trained with the proposed KAL loss against the standard MSE loss. All methods are evaluated on the same object set with an equal number of candidates. Following the evaluation protocol of DRO, we report two complementary metrics:
@@
Result Analysis. Table I summarizes performance across ShadowHand, Allegro, and Barrett. MorphArtGrasp achieves the highest overall success rate (91.9%) while also being the most efficient ($0.353\,\text{s}$). Compared to DRO, we obtain substantial gains on ShadowHand (+10.7%) and Allegro (+1.1%). The only exception is the Barrett hand, where DRO obtains a slightly higher success rate. This can be attributed to DRO’s optimization-based nature, which is more effective on hands with lower degrees of freedom, as also suggested by the trend in Table I.
More notably, MorphArtGrasp outperforms DexGraspNet across all hand types and achieves significantly better efficiency. We found DexGraspNet highly sensitive to hyperparameter settings across different embodiments. Despite extensive tuning, performance varied, highlighting a key drawback of optimization-based methods-reliance on hand-specific hyperparameter tuning. In contrast, MorphArtGrasp maintains consistently high performance across all hand types without requiring hand-specific parameter adjustment.
@@
TABLE II: Grasp success rates and improvement (%) using wrist poses from baseline methods with MorphArtGrasp.
@@ IV-D Articulation Prediction with Baseline Wrist Poses
We further evaluate MorphArtGrasp on the unseen-object test set using wrist poses generated by existing baselines (Sec. IV-C). Specifically, we take the wrist poses proposed by DFC [9], DexGraspNet [20], and DRO [22], and apply MorphArtGrasp to generate the corresponding articulations. This experiment examines whether the predicted articulations can improve grasp success when paired with wrist poses from external methods (Q2).
Table II summarizes the grasp success rates. The numbers in brackets indicate the improvement in success rate compared to the baseline methods. The results indicate that MorphArtGrasp demonstrates consistent improvements when applied to baseline wrist poses across most configurations, suggesting the potential for greater performance increase when paired with higher quality wrist pose generation.
Discussion. While in the DRO setup, MorphArtGrasp improves grasp success for ShadowHand (+14.7%) and Allegro (+5.3%), we observe a 6.3% drop for the Barrett hand. Visualization indicates that DRO-generated Barrett grasps typically succeed through form-closure, achieved when the wrist pose is close to the object. In contrast, MorphArtGrasp, guided by the Kinematic-Aware Articulation Loss (KAL), emphasizes fingertip-relevant motions and thus favors wrist poses farther from the object, enabling force-closure grasps (see Fig. 5). This does not reflect a drawback of KAL itself, but rather a shift in preferred grasping strategy. For hands with limited DoFs such as Barrett, this shift may occasionally reduce compatibility with DRO’s wrist pose proposals, whereas for more dexterous hands (ShadowHand, Allegro) it leads to notable gains.
Figure 5: Comparison of grasping strategies on the Barrett hand. Left: a form-closure grasp favored by DRO, where the wrist pose is close to the object and fingers wrap around it. Right: a force-closure grasp favored by MorphArtGrasp with KAL, where fingertip contacts dominate.
@@ IV-E Few-Shot Generalization to Unseen Hand
To evaluate the ability of MorphArtGrasp to generalize across embodiments, we conduct a few-shot adaptation experiment on an unseen hand: the Robotiq 3-Finger (11 DoF), addressing Q3. We adopt the Robotiq3F dataset from MultiGripperGrasp dataset [2] and randomly sample 100 objects, each with 10 grasp poses, as the training set for this new hand. The model is then fine-tuned with these limited examples and evaluated in Isaac Gym using the predicted wrist poses (following the same mapping procedure as in Sec. IV-C).
For evaluation, we use our unseen object test set. The simulation results show that MorphArtGrasp achieves a grasp success rate of 85.6%, demonstrating strong few-shot generalization to previously unseen hands.
@@ IV-F Real-Robot Experiments
To further validate MorphArtGrasp in the physical world (Q4), we deploy the Robotiq 3-Finger hand mounted on a Franka Panda arm. Object point clouds are captured using three Intel RealSense D435 cameras placed around the workspace. For each test object, wrist poses are predicted by 6-DoF GraspNet and mapped to the hand-specific wrist frame, after which MorphArtGrasp (fine-tuned on the Robotiq 3F in Sec. IV-E) predicts the articulation. The wrist pose and articulation are then executed to determine grasp validity.
We evaluate grasp performance on a set of 10 previously unseen objects, with 10 grasp attempts per object. Candidate poses are filtered based on Robotiq 3-Finger DoF constraints and physical feasibility (e.g., table collisions). Example predicted grasps on the test objects are illustrated in Fig. 4. The system achieves a grasp success rate of 87%, demonstrating that MorphArtGrasp transfers effectively from simulation to the real world.
@@ V CONCLUSION
We presented MorphArtGrasp, an eigengrasp-based framework for cross-embodiment dexterous grasp generation. MorphArtGrasp converts each hand’s URDF into structured morphological tokens to obtain a morphology embedding and hand-specific eigengrasps, and supervises articulation prediction with a Kinematic-Aware Articulation Loss that captures morphology-dependent kinematics beyond joint-space errors. Extensive experiments in simulation and on real hardware demonstrate strong generalization to novel objects and effective transfer across multiple dexterous hands.

### v3

# MachaGrasp: Morphology-Aware Cross-Embodiment Dexterous Hand Articulation Generation for Grasping

@@ Abstract
Dexterous grasping with multi-fingered hands remains challenging due to high-dimensional articulations and the cost of optimization-based pipelines. Existing end-to-end methods require training on large-scale datasets for specific hands, limiting their ability to generalize across different embodiments. We propose MachaGrasp, an eigengrasp-based, end-to-end framework for cross-embodiment grasp generation. From a hand’s morphology description, we derive a morphology embedding and an eigengrasp set. Conditioned on these, together with the object point cloud and wrist pose, an amplitude predictor regresses articulation coefficients in a low-dimensional space, which are decoded into full joint articulations. Articulation learning is supervised with a Kinematic-Aware Articulation Loss (KAL) that emphasizes fingertip-relevant motions and injects morphology-specific structure. In simulation on unseen objects across three dexterous hands, MachaGrasp attains a 91.9% average grasp success rate with $\mathord{<}0.4\,\mathrm{s}$ inference per grasp. With few-shot adaptation to an unseen hand, it achieves 85.6% success on unseen objects in simulation, and real-world experiments on this few-shot-generalized hand achieve an 87% success rate. The code and additional materials are available on our project website https://connor-zh.github.io/MachaGrasp/.
@@ I INTRODUCTION
In this paper, we propose MachaGrasp, a framework for cross-embodiment dexterous grasp generation. MachaGrasp extracts eigengrasps and morphology embeddings directly from a robotic hand’s Unified Robot Description Format (URDF), explicitly incorporating kinematic and geometric constraints. Given an object point cloud and a target wrist pose, an amplitude predictor estimates articulation coefficients in the eigengrasp space, which are then decoded into full joint configurations. To supervise training, we introduce a Kinematic-Aware Articulation Loss (KAL) that emphasizes fingertip-relevant motions and implicitly encodes morphology-specific information, in contrast to naïve per-joint regression objectives (e.g., mean squared error) that overemphasize raw joint deviations.
We demonstrate the effectiveness of the proposed framework through extensive evaluations in both simulation and real-world settings. In simulation on unseen objects across three dexterous hands, MachaGrasp achieves a 91.9% average grasp success rate with $0.4\,\text{s}$ inference per grasp; with few-shot adaptation to an unseen hand it attains 85.6% success on unseen objects in simulation, and real-world experiments on the same few-shot–generalized hand achieve 87% success.
@@
  • We propose MachaGrasp, an eigengrasp-based framework for cross-embodiment dexterous grasp generation that predicts articulations in an end-to-end manner.
@@ II-B Cross-Embodiment Dexterous Grasping
Another line of research focuses on enabling dexterous grasping across different robotic hands by leveraging hand-agnostic representations or shared action spaces. Cross-embodiment reinforcement learning approaches, such as the eigengrasp-based method by Yuan et al. [24], employ a shared low-dimensional action space to allow a single policy to operate across multiple hands, although training is often sample-intensive. More recently, GET-Zero [12] introduces a graph-based embodiment transformer that conditions policy learning on the structural graph of the hand, achieving zero-shot transfer to unseen morphologies in dexterous in-hand manipulation tasks. While highly effective for manipulation, its focus is on policy learning rather than generating feasible grasps, leaving a gap that MachaGrasp addresses by focusing on effective grasp generation for diverse hand morphologies.
@@
Figure 2: Method Overview. MachaGrasp processes hand URDF, object point cloud, and wrist pose through specialized encoders to generate morphology embedding $\boldsymbol{m}$, point cloud embedding $\boldsymbol{p}$, and pose encoding $\boldsymbol{h}$. These are fused to predict eigengrasp amplitudes $\boldsymbol{a}$, which are combined with eigengrasps $E$ to produce the final articulation vector $\boldsymbol{q}$.
@@ III-B Method Overview
An overview of MachaGrasp is shown in Fig. 2. The framework takes as input the hand URDF, the object point cloud, and the wrist pose. The hand URDF is first processed by the morphology encoder (Sec. III-C), which produces a morphology embedding ($\boldsymbol{m}$) together with a set of eigengrasps ($E$). The object point cloud is then processed by the object encoder (Sec. III-D) to extract a point cloud embedding ($\boldsymbol{p}$). These embeddings, combined with the wrist pose ($\boldsymbol{h}$), are fused to form conditioned eigengrasp tokens. The amplitude predictor (Sec. III-E) takes these tokens as input and estimates the amplitudes for each eigengrasp ($a$). Finally, the articulation vector ($\boldsymbol{q}$) is obtained by the dot product between the predicted amplitudes and the eigengrasps, and together with the wrist pose, defines the final hand pose.
@@ III-C Morphology Encoder
A central component of MachaGrasp is the morphology encoder (illustrated in Fig. 3), which learns structured representations of dexterous hands directly from their URDF descriptions. In contrast to prior work that encodes hand morphology solely as mesh point clouds [15, 9, 20], MachaGrasp leverages the URDF to extract explicit kinematic constraints and geometric primitives. Based on this structured information, the encoder predicts both the eigengrasps of the hand and a compact latent representation that captures its underlying morphology.
@@ IV EXPERIMENTS
  • Q1: How well does MachaGrasp perform in realistic settings where wrist poses are predicted by a naive pose predictor, compared to state-of-the-art baselines?
  • Q2: Does integrating articulation predictions from MachaGrasp with the wrist poses generated by existing baselines (GraspIt, DexGraspNet, DRO) improve their grasp success rates?
  • Q3: Can MachaGrasp generalize to previously unseen robotic hands with only few-shot adaptation?
@@ IV-A Datasets
We train MachaGrasp on three dexterous robotic hands: ShadowHand (5 fingers, 22 DoF), Allegro Hand (4 fingers, 16 DoF), and Barrett Hand (3 fingers, 8 DoF). For Allegro and Barrett, we adopt grasp data from the MultiGripperGrasp dataset [2]. Since the ShadowHand subset in MultiGripperGrasp suffers from severe self-collision artifacts, we instead use DexGraspNet [20] for ShadowHand data.
@@ IV-C Articulation Prediction with Predicted Wrist Pose
To answer Q1, we evaluate MachaGrasp in a realistic setting where wrist poses are predicted rather than provided. Specifically, we use 6-DoF GraspNet [11] to generate parallel-gripper candidates per object, and map each candidate to a hand-specific wrist pose using a calibrated rigid transform $\Delta_{\text{hand}}$:
@@
For each retained wrist pose $T_{\text{wrist}}^{(\text{hand})}$, MachaGrasp predicts an articulation, and the resulting grasp is evaluated in simulation. Experiments are conducted on a test set of 28 unseen objects, with 50 candidate grasps per object. Example grasps generated by MachaGrasp are visualized in Fig. 1.
@@
As an ablation, we also compare MachaGrasp trained with the proposed KAL loss against the standard MSE loss. All methods are evaluated on the same object set with an equal number of candidates. Following the evaluation protocol of DRO, we report two complementary metrics:
@@
Result Analysis. Table I summarizes performance across ShadowHand, Allegro, and Barrett. MachaGrasp achieves the highest overall success rate (91.9%) while also being the most efficient ($0.353\,\text{s}$). Compared to DRO, we obtain substantial gains on ShadowHand (+10.7%) and Allegro (+1.1%). The only exception is the Barrett hand, where DRO obtains a slightly higher success rate. This can be attributed to DRO’s optimization-based nature, which is more effective on hands with lower degrees of freedom, as also suggested by the trend in Table I.
More notably, MachaGrasp outperforms DexGraspNet across all hand types and achieves significantly better efficiency. We found DexGraspNet highly sensitive to hyperparameter settings across different embodiments. Despite extensive tuning, performance varied, highlighting a key drawback of optimization-based methods-reliance on hand-specific hyperparameter tuning. In contrast, MachaGrasp maintains consistently high performance across all hand types without requiring hand-specific parameter adjustment.
@@
TABLE II: Grasp success rates and improvement (%) using wrist poses from baseline methods with MachaGrasp.
@@ IV-D Articulation Prediction with Baseline Wrist Poses
We further evaluate MachaGrasp on the unseen-object test set using wrist poses generated by existing baselines (Sec. IV-C). Specifically, we take the wrist poses proposed by DFC [9], DexGraspNet [20], and DRO [22], and apply MachaGrasp to generate the corresponding articulations. This experiment examines whether the predicted articulations can improve grasp success when paired with wrist poses from external methods (Q2).
Table II summarizes the grasp success rates. The numbers in brackets indicate the improvement in success rate compared to the baseline methods. The results indicate that MachaGrasp demonstrates consistent improvements when applied to baseline wrist poses across most configurations, suggesting the potential for greater performance increase when paired with higher quality wrist pose generation.
Discussion. While in the DRO setup, MachaGrasp improves grasp success for ShadowHand (+14.7%) and Allegro (+5.3%), we observe a 6.3% drop for the Barrett hand. Visualization indicates that DRO-generated Barrett grasps typically succeed through form-closure, achieved when the wrist pose is close to the object. In contrast, MachaGrasp, guided by the Kinematic-Aware Articulation Loss (KAL), emphasizes fingertip-relevant motions and thus favors wrist poses farther from the object, enabling force-closure grasps (see Fig. 5). This does not reflect a drawback of KAL itself, but rather a shift in preferred grasping strategy. For hands with limited DoFs such as Barrett, this shift may occasionally reduce compatibility with DRO’s wrist pose proposals, whereas for more dexterous hands (ShadowHand, Allegro) it leads to notable gains.
Figure 5: Comparison of grasping strategies on the Barrett hand. Left: a form-closure grasp favored by DRO, where the wrist pose is close to the object and fingers wrap around it. Right: a force-closure grasp favored by MachaGrasp with KAL, where fingertip contacts dominate.
@@ IV-E Few-Shot Generalization to Unseen Hand
To evaluate the ability of MachaGrasp to generalize across embodiments, we conduct a few-shot adaptation experiment on an unseen hand: the Robotiq 3-Finger (11 DoF), addressing Q3. We adopt the Robotiq3F dataset from MultiGripperGrasp dataset [2] and randomly sample 100 objects, each with 10 grasp poses, as the training set for this new hand. The model is then fine-tuned with these limited examples and evaluated in Isaac Gym using the predicted wrist poses (following the same mapping procedure as in Sec. IV-C).
For evaluation, we use our unseen object test set. The simulation results show that MachaGrasp achieves a grasp success rate of 85.6%, demonstrating strong few-shot generalization to previously unseen hands.
@@ IV-F Real-Robot Experiments
To further validate MachaGrasp in the physical world (Q4), we deploy the Robotiq 3-Finger hand mounted on a Franka Panda arm. Object point clouds are captured using three Intel RealSense D435 cameras placed around the workspace. For each test object, wrist poses are predicted by 6-DoF GraspNet and mapped to the hand-specific wrist frame, after which MachaGrasp (fine-tuned on the Robotiq 3F in Sec. IV-E) predicts the articulation. The wrist pose and articulation are then executed to determine grasp validity.
We evaluate grasp performance on a set of 10 previously unseen objects, with 10 grasp attempts per object. Candidate poses are filtered based on Robotiq 3-Finger DoF constraints and physical feasibility (e.g., table collisions). Example predicted grasps on the test objects are illustrated in Fig. 4. The system achieves a grasp success rate of 87%, demonstrating that MachaGrasp transfers effectively from simulation to the real world.
@@ V CONCLUSION
We presented MachaGrasp, an eigengrasp-based framework for cross-embodiment dexterous grasp generation. MachaGrasp converts each hand’s URDF into structured morphological tokens to obtain a morphology embedding and hand-specific eigengrasps, and supervises articulation prediction with a Kinematic-Aware Articulation Loss that captures morphology-dependent kinematics beyond joint-space errors. Extensive experiments in simulation and on real hardware demonstrate strong generalization to novel objects and effective transfer across multiple dexterous hands.